OpenStudy (anonymous):

Sara invested $8,000, part at 6% and the rest at 9% interest per year. How much did she invest at each rate if her total return per year on the investment was $660?

OpenStudy (mathstudent55):

I = Prt I = interest, P = principal, r = rate, t = time Let's say she invested x at 6%. That means she invested 8000 - x at 9%. Can you write an expression for the interest earned at 6% and an expression for the interest earned at 9%?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

is it .06x+.09y=660?

OpenStudy (mathstudent55):

I = Prt I = I P = x r = 6% t = 1 Use the equation above and the info above to write one expression for the interest earned at 6%

OpenStudy (mathstudent55):

Yes, but you can do one thing now.

OpenStudy (mathstudent55):

You used x for the amount invested at 6% and y for the amount invested at 9%. What do we know about x and y?

OpenStudy (anonymous):

they're different variables?

OpenStudy (mathstudent55):

Where does the 8000 come in?

OpenStudy (anonymous):

.06

OpenStudy (anonymous):

and .09?

OpenStudy (anonymous):

so i plug in 800 for x and y?

OpenStudy (mathstudent55):

No. Your equation above is correct. You used x as the part of the $8000 amount invested at 6%. You used y as the part of the $8000 amount invested at 9%. Using x, y, and 8000, can you write the relation between x and y?

OpenStudy (mathstudent55):

The total amount invested is $8000. x was invested at 6%. y was invested at 9%. x + y = ?

OpenStudy (anonymous):

.06(800)+.09y=660?

OpenStudy (anonymous):

then y =6800

OpenStudy (anonymous):

but i don't understand what to do next?

OpenStudy (mathstudent55):

No. You need a second equation which is x + y = 8000 Now you have a system of equations: 0.06x + 0.09y = 660 x + y = 8000

OpenStudy (anonymous):

.06x+.09(6800)=660 x=-800? is that right?

OpenStudy (anonymous):

so for the second eqatuion do i plug in x=-800 and y=6800?

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

can you just give me the answer please?

OpenStudy (mathstudent55):

Where do you get 800 from. It's 8000. Ok, we have two equations, 0.06x + 0.09y = 660 x + y = 8000 Let's use the substitution method to solve the system of equations.

OpenStudy (mathstudent55):

We solve x + y = 8000 for y: y = 8000 - x Now we replace y with 8000 - x in the first equation: 0.06x + 0.09(8000 - x) = 660 0.06x + 720 - 0.09x = 660 -0.03x = -60 x = 2000 x + y = 8000 2000 + y = 8000 y = 6000 Answer: $2000 invested at 6% and $6000 invested at 9%

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!